A photon has a frequency of 8.99 X 10^8 Hz. What is the energy of this photon? (h= 6.626 X 10^-34 J X S)
1
Expert's answer
2020-04-27T01:52:48-0400

The energy will be calculated according to the equation:

"E = h•\nu" = 6.626•10-34(J•s) • 8.99•108(Hz = s-1) = 59.6•10-26 J

Answer: E=59.6•10-26 J